# EdgeAlpha: Distributed Process Discovery at

### v3

@@ Abstract
Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT) in a smart home, factory or hospital setting. The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs. Further, privacy considerations often prevent data collection at a central location in the first place. To address this challenge, this paper introduces EdgeAlpha, an algorithm for distributed process discovery operating directly on sensor nodes on a stream of real-time event data. Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model. Our analysis shows that even with over 600 activities, like in the Hospital Log, the average number of predecessors per activity remains below 7. This enables efficient determination of predecessors for each event, reducing the communication overhead by up to 96% compared to querying all nodes and enhancing scalability. Additionally, our analysis demonstrates that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups of 40 reduces the average queried nodes per event to less than 2.5%. EdgeAlpha ensures that the resulting process model is identical to the original Alpha Miner.
Index Terms:
@@
Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability
I Introduction
In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from Internet of Things (IoT) sensors, for example, in a smart manufacturing, smart home or smart health setting [1, 2, 3, 4]. However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [5, 6, 7]. This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [8]. Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place. For example, patient data in hospitals, personal data of workers, data in smart home setups, or company secrets in cross-organizational processes are commonly not shared for analysis [9, 10, 11]. We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
@@ I Introduction
Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
In this paper, we introduce EdgeAlpha, an algorithm for distributed process mining based on the Alpha Miner [12]. Utilizing the constrained compute and storage resources of sensor nodes, further referred to as nodes, we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location. In contrast to Gatta et al. [9] which enable cross-organizational process mining through the transmission of data aggregates, EdgeAlpha distributes the computation of the footprint matrices (FMs) to the sensor-node level. This approach allows the collection of aggregated data within an organization that would otherwise not be obtainable in raw form due to, for example, too high data velocity or data protection regulations.
@@
For each event that a node in EdgeAlpha detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig. 1. Thus, over time, each node constructs a partial FM that represents the view of this node on the process. To derive the complete process model, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix. EdgeAlpha enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with a central entity and (b) privacy-preserving process mining, as nodes only store their own as well as predecessor and successor events.
@@
To determine an event’s predecessor event, the node that records this particular event potentially has to query all other nodes, which can lead to large communication overheads. In this paper, we show that activities of real-world processes statistically only have a few predecessors activities, which further only rarely change throughout the lifetime of a process. Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [13] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. As a result, for each event, a node can very efficiently determine the predecessor reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events. Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes. In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor and a baseline that immediately queries the correct node. Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner and show that in four out of five datasets an intermediate result FM already complies with the FM of the Alpha Miner by more than 90% after less than 5% of the events of the datasets.
EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [14].
@@
  • We introduce EdgeAlpha, a distributed Alpha Miner operating directly where the events originate.
  • We analytically and experimentally show the correctness of the algorithm.
  • We demonstrate that EdgeAlpha’s intermediate results quickly converge to the final result of the Alpha Miner.
  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach and by 97.5% when batching queries with a batch size of 40.
  • We ensure that the nodes only share aggregated event data with a central entity.
  • We implement EdgeAlpha and publish its code open-source [^1] Available at: redacted..
The remainder of the paper is structured as follows. Section II introduces the necessary background. Next, Section III presents related work on distributed process discovery. Afterwards, we introduce the design of EdgeAlpha in Section IV. In Section V, we evaluate EdgeAlpha. We conclude the paper in Section VI.
II Preliminaries and Background
In this section, we provide the required preliminaries and background. We provide definitions of direct successions and footprint matrices and recapitulate the steps of the Alpha Miner Algorithm [12] that are relevant to our approach.
We refer to Peterson [15] for details on Petri nets and use standard definitions for activities, event logs, cases, and traces [16, 17]. In the following, let $L$ be an event log over the activity set $T$. We define direct successions and the resulting relationships between two activities following the definitions by Gatta et al. [9].
Definition 1 (Direct Succession)
@@ II Preliminaries and Background
Definition 2 (Causality, No Direct Succession)
With the help of Definition 1 we now define causality (1) and no direct succession (2) between two activities.
$a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (1)
$a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (2)
Next, we recapitulate the initial steps of the Alpha Miner before later in the paper introducing a distributed variant. First, the algorithm identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$. Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively. Subsequently, the algorithm determines all direct successions in $L$ which can be stored in an FM. This FM is then the basis for the further calculations in the algorithm and the resulting Petri net.
A FM stores information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is an element of $\{0,\dots,|T_{L}|-1\}$. This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
$FM_{ij}=1\quad$ $\Leftrightarrow\quad i>_{L}j$
@@
III Related Work
@@
Van der Aalst [18] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. Here, models rather than FMs are created in a distributed manner. Evermann and Assadipour [19] leverage Map-Reduce [20] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes. However, their approach still relies on an event log that is initially stored at a central location. Subsequent work [21] extends their approach to the Flexible Heuristics Miner [22] and evaluates the performance of the Map-Reduce-based Alpha Miner and the Flexible Heuristics Miner. While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only storing data aggregates at a central location.
Gatta et al. [9] address privacy concerns in the healthcare domain: In their work, each hospital center aggregates its own event data into an FM, such that the others only see the aggregated versions of the data. Like other works [18, 19, 21], they subsequently merge the FMs. In contrast, EdgeAlpha distributes the FM computations to the sensor nodes that record the events, enabling data aggregation at an earlier stage. This allows organizations to collect and utilize data in aggregated form that would otherwise be restricted in its raw form due to data protection regulations. As a result of processing events directly at their sources, EdgeAlpha increases scalability and reduces privacy risks.
IV Design
In this section, we first discuss our assumptions before detailing on the design of EdgeAlpha, introducing our strategy to find predecessors and proving correctness.
IV-A Assumptions on Events and Event Logs
In adherence to common practice, we assume events with the attributes case ID, activity name, and timestamp, whereas the case ID is unique for each case. Our evaluation shows these assumptions are both realistic and applicable in numerous real-world scenarios. For example, in a hospital each patient is commonly assigned a fixed ID that can serve as a case ID and the events such as blood tests or ultrasound are predefined activities. Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log. Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log. As we operate on distributed event logs, we assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
IV-B EdgeAlpha Setting
@@
First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing the EdgeAlpha algorithm. We employ EdgeAlpha in typical IoT application settings: A network of sensor nodes that are, for example, integrated into machines in a smart factory or smart health setting. These sensors can communicate with each other, either directly or via the Internet. Further, we assume that a membership management allows nodes to gain knowledge of other nodes and how to reach them, as common in IoT applications. Together, these allow a node to find the processor and successor of an event. Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP [23]. This is essential, as accurate timestamps ensure proper event ordering. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node detects exactly one type of activity. Note that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
IV-C EdgeAlpha Algorithm: Overview
@@
EdgeAlpha consists of two phases: Phase 1 is initiated upon event triggering, and Phase 2 is activated upon a process model request. In Phase 1, nodes communicate to determine predecessors when recording an event and note their directly-follows relations in their local FMs. As a result, every node continuously updates its local, partial FM. Other than Phase 1, Phase 2 is activated on demand whenever an entity requests a process model. Then, all nodes send their partial FMs to a central location, which constructs an FM. We define both phases, starting with the basic version of Phase 1, which sends a message to all nodes in the system for each event, followed by a more efficient variant and event batching.
IV-D Phase 1 – Event Ordering & Partial Footprint Matrices

### v4

# EdgeMiner: Distributed Process Mining at the Data Sources

@@ Abstract.
Process mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT). The volume and velocity of data generated by such sensors makes it increasingly challenging to efficiently process the data by traditional process discovery algorithms, which operate on a centralized event log. This paper presents EdgeMiner, an algorithm for distributed process mining operating directly on sensor nodes on a stream of real-time event data. In contrast to centralized algorithms, EdgeMiner tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. As EdgeMiner aggregates direct successions on the individual nodes, the raw data does not need to be stored centrally, thus improving both scalability and privacy. We analytically and experimentally show the correctness of EdgeMiner. In addition, our evaluation results show that EdgeMiner determines predecessors for each event efficiently, reducing the communication overhead by up to 96% compared to querying all nodes. Further, we show that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups reduces the average queried nodes per event to less than 2.5%.
@@
Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability, Internet of Things
1. Introduction
Mining the never-ending data stream of events sensed by the Internet of Things (IoT) provides deep insights into, for example, processes in smart manufacturing, smart homes, and healthcare and is known as process mining (Bertrand et al., 2022; Fernandez-Llatas et al., 2015; Malburg et al., 2022; Janssen et al., 2020). However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud (Ge et al., 2018; Zaslavsky et al., 2013; van der Aalst, 2013). This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities (van der Aalst, 2016). Parsing such event logs, established algorithms commonly construct so-called footprint matrices (FMs), in which they store the order of events, so-called directly follows relations (van der Aalst et al., 2004; Leemans et al., 2013; Weijters et al., 2006). Next to the scalability challenge, such collection of events at a central location causes privacy concerns (Gatta et al., 2019; Mannhardt et al., 2019; Pika et al., 2020).
@@ 1. Introduction
Figure 1. While traditional process mining (left) collects all events at a central entity, EdgeMiner (right) processes them directly at the source, and only exchanges aggregates (partial footprint matrixes), increasing scalability and privacy.
[…]
[From centralized to distributed process mining.]While classical process mining (left) collects all events on a central entity, EdgeMiner processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
@@
In this paper, we introduce EdgeMiner, an algorithm for distributed process mining: it composes FMs, the key building-block of process mining, in a distributed manner within a network of sensor nodes. For each event that a node in EdgeMiner detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig. 1. Thus, over time, each node constructs a partial FM that represents the view of this node on the direct successions of this process. To derive the complete FM of the process, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix. As a result, EdgeMiner enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with the requesting entity and (b) privacy-preserving process mining, as nodes locally merely store their own as well as predecessor and successor events and a requesting entity only receives aggregate FMs instead of raw event streams. We prove, in this paper, that EdgeMiner guarantees an FM identical to the FM when using a centralized approach.
@@
Analyzing real-world datasets, our evaluation shows that activities statistically only have a few predecessor activities, which further only rarely change throughout the lifetime of a process. For example, even in the large Hospital Log (van Dongen, 2011), comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. As a result, for each event, a node can very efficiently determine the predecessor, reducing the communication overhead by up to 96% compared to the naive baseline of querying all nodes in the network. Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeMiner can reduce the average number of queried nodes per event to less than 2.5% of all nodes. As a final result, we show that for most datasets, an intermediate FM already complies with the final FM by more than 90% after less than 5% of the events of the dataset.
@@
  • We introduce EdgeMiner, a distributed process mining algorithm enabling distributed process discovery and conformance checking by operating directly where the events originate.
  • We ensure that the nodes only share aggregated event data with a requesting entity.
  • We analytically and experimentally show the correctness of the merge of EdgeMiner.
  • We demonstrate that EdgeMiner’s interim results quickly converge to the footprint matrix of the whole event log.
  • We show that EdgeMiner reduces the message count for the predecessor determination by up to 96% compared to the naive approach and, for example, by 97.5% when batching queries with a batch size of 40.
  • We implement EdgeMiner and publish its code open-source [^1] Available at: https://github.com/ds-kiel/EdgeMiner.
The remainder of the paper is structured as follows. Section 2 introduces the necessary background. Next, Section 3 presents related work on distributed process mining. Afterwards, we introduce the design of EdgeMiner in Section 4. In Section 5, we evaluate EdgeMiner. We conclude the paper in Section 6.
2. Preliminaries and Background
In this section, we give a brief introduction into process mining and provide definitions of events, traces, event logs, direct successions and footprint matrices. In process mining, we commonly discover a process model from an event log (process discovery) or check that a process follows a given model (conformance checking) (van der Aalst, 2011; van der Aalst, 2016; van der Aalst et al., 2012; van der Aalst, 2012). A common way to represent a process model is a Petri net (Peterson, 1977; Murata, 1989).
Events and Event Logs: Process mining algorithms operate on event logs. Each event in the log typically contains at least three attributes: an activity, a timestamp, and a case ID. The case ID is used to distinguish different process instances. A trace is a sequence of events corresponding to a particular case ordered by timestamp.
Direct Successions: In the following, let $L$ be an event log over the activity set $T$. We define direct successions and therewith definable relationships between two activities, following the definition by Gatta et al. (Gatta et al., 2019).
Definition 2.1 (Direct Succession).
@@ 2. Preliminaries and Background
Definition 2.2 (Causality, No Direct Succession).
With the help of Definition 2.1 we now define causality (1) and no direct succession (2) between two activities.
(1) $a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$
(2) $a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$
Footprint Matrix: A footprint matrix (FM) stores information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is an element of $\{0,\dots,|T_{L}|-1\}$. An entry denotes how often the activity corresponding to the row was followed by the activity of the column. This way, $FM_{ij}\in\mathbb{N}_{0}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
$FM_{ij}>0\quad$ $\Leftrightarrow\quad i>_{L}j$
@@
Alpha Miner: To illustrate process mining and its use of FMs, we recapitulate the main steps of the seminal process discovery algorithm Alpha Miner (Algorithm 1). Later, we discuss how newer discovery algorithms extend the concept of the alpha miner and its FMs.
First, the algorithm denotes all activities occurring in the event log $L$ as $T_{L}$. Each activity of this set corresponds to a transition in the final output model computed by the algorithm. Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively. Subsequently, the algorithm determines all direct successions in $L$ which it stores in an FM. This FM is the basis for the further calculations in the algorithm and the resulting Petri net. Next, the algorithm generates a set $X_{L}$, which comprises pairs of activity sets. Each element of the first set is directly followed by every element of the second set, although not vice versa. Moreover, activities within each set do not directly succeed one another. The Alpha Miner minimizes the set of pairs, followed by the derivation of Petri net places from the minimized pairs in the next step. Finally, the algorithm adds arcs between these places and transitions, forming the resulting Petri net.
Algorithm 1 Alpha Miner
1:$T_{L}=\{t\in T|\exists\sigma\in L:t\in\sigma\}$
2:$T_{I}=\{first(\sigma)\in T|\forall\sigma\in L\}$
3:$T_{O}=\{last(\sigma)\in T|\forall\sigma\in L\}$
[…]
6:$P_{L}=\{p_{(A,B)}|(A,B)\in Y_{L}\}\cup\{i_{L},o_{L}\}$
[…]
8:$\alpha(L)=(P_{L},T_{L},F_{L})$
To form the set $X_{L}$, we need to have knowledge of the $\rightarrow_{L}$ and #$\#_{L}$ relationships between activities. These relations solely depend on the direct successions between the activities. An FM stores exactly those direct successions. For the Alpha Miner, it is not necessary to know the number of direct successions, just whether they exist or not.
Heuristics Miner and Inductive Miner: Newer process discovery algorithms, extent on the concept of the Alpha Miner while keeping the FM as a key building block. For example, the Heuristics Miner starts with the construction of a dependency graph (Weijters et al., 2006). To build the dependency graph, it calculates a so-called dependency measure between each two activities, depending on the number of direct successions between the activities. As an FM stores all direct successions, the miner use it for this calculation. The Inductive Miner also relies on direct successions stored in an FM. The core concept is to construct a directly-follows graph, which visually represents the FM along with start and end activity sets. From this graph, a process tree is formed by repeatedly partitioning the nodes of the directly-follows graph, capturing the structure of the process step by step.
3. Related Work
@@
Today’s approaches to distributed process discovery focus primarily on multiple compute nodes to speed up the processing of ever-growing, centralized event logs: Van der Aalst (van der Aalst, 2013) presents one of the first methods for distributed process discovery. He introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. In contrast to EdgeMiner, he creates models rather than FMs in a distributed manner. Others (Evermann and Assadipour, 2014) leverage Map-Reduce (Dean and Ghemawat, 2008) to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes to handle the ever-growing size of event logs. However, their approach still relies on an event log that is initially stored at a central location. Later works  (Evermann, 2016) bring Map-Reduce to the Flexible Heuristics Miner (Weijters and Ribeiro, 2011). Gatta et al. (Gatta et al., 2019) extends the primary motivation of the previous work (i.e., performance and scalability) to include privacy concerns in the healthcare domain, but still focuses on one event log per hospital. In contrast to all these, EdgeMiner emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only sending aggregate information, i.e., partial FMs, to the requesting node.
In the context of conference checking, the Single-Entry Single-Exit (SESE) breaks down large process models and event logs into smaller sub-processes to analyze them independently (Munoz-Gama et al., 2014). Similarly, passage-based decomposition (van der Aalst and Verbeek, 2014) and Projected Conformance Checking (PCC) (Leemans et al., 2018) split logs into sets of activities that share direct successor-predecessor relationships: By solving each set independently and merging the results, it distributes, for example, conformance checking across multiple nodes. Although these approaches efficiently handle large and complex event logs, they remain centralized. One recent work presents conformance checking in a distributed manner also based on footprint-matrices, yet – in contrast to EdgeMiner– their work does not cover the underlying mechanism to distribute the calculation of the matrices (Andersen et al., 2024).
Streaming process mining (Burattin, 2022) takes a different approach to distributed process mining. It focuses mainly on the real-time analysis of streams of events (i.e., event streams). Generally, those algorithms maintain computational state according to on previous events (Burattin et al., 2014; Hassani et al., 2015; van Zelst and et. al, 2019; Burattin et al., 2012) and employ approximations such as lossy counting (Burattin et al., 2014). While those methods are also motivated by scalability and performance, they are not designed to compute FMs at the sensor nodes.
In contrast, EdgeMiner distributes the FM computation and each sensor node computes a partial FM, enabling data aggregation at an earlier stage. As a result of processing events directly at their sources, EdgeMiner increases scalability, reduces privacy risks and enables both process discovery and online conformance checking.
4. Design
In this section, we begin by discussing our assumptions and setting. Next, we give a design overview of EdgeMiner, detail on its design, introduce optimizations, prove correctness, and discuss how EdgeMiner enables distribution of well-known process mining algorithms.
4.1. Assumptions and Setting
4.1.1. Assumptions on Events and Event Logs
In adherence to common practice, we assume events with the attributes case ID, activity name, and timestamp, whereas the case ID is unique for each case. We assume timestamps within the same case to be distinct, ensuring unique FMs. In our evaluation in Section 5, we show that our assumptions are both realistic and applicable in numerous real-world scenarios.
4.1.2. EdgeMiner Setting
@@
EdgeMiner targets typical IoT application settings: A network of IoT nodes that are, for example, integrated into machines in a smart factory or smart health setting. These sensors can communicate with each other, either directly or via the Internet. Further, we assume that a membership management allows nodes to gain knowledge of other nodes and how to reach them, as common in IoT applications. Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP (Mills, 1985). This is essential, as accurate timestamps ensure proper event ordering. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node detects exactly one type of activity. Note that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
4.2. EdgeMiner Algorithm
@@
EdgeMiner consists of two phases: Phase 1 operates continuously upon each event detected, and Phase 2 is executed upon a footprint matrix request. In Phase 1, nodes communicate to determine predecessors when recording an event and note their directly-follows relations in their local FMs. As a result, every node continuously updates its local, partial FM. Other than Phase 1, Phase 2 runs on demand, whenever a (central) entity requests the FM. Then, all nodes send their partial FMs to the requesting entity, which constructs the FM.
4.2.1. Phase 1 – Event Ordering & Partial Footprint Matrices